# Blood Flow in Elastic Vessel Using PINNH. Zhang, R. Chan and X.C. Tai

### v1

@@ 1 Introduction
To address these challenges, we propose a neural solver for the Navier-Stokes equation in deformable vessels using Physics-Informed Neural Networks (PINNs)[27], which has achieved some success on flow problems in non-moving domains [17, 8]. In our methods, we take some networks to produce displacement, pressure and velocity of our problem. To acquire those networks, we first convert the partial differential equations of the formulated physical model into residual forms and use them as loss functions. Next, the automatic differentiation is used to represent all the differential operators. Then, we train the networks to minimize those losses. With those losses getting approximately zero, the networks can be an approximation of the functions for displacement, pressure, and velocity that are used to solve the modeled problem.
This method is a mesh-free approach, eliminating the need for discretization and meshing for the computational domain. This feature makes it highly efficient and straightforward to add other objects into the geometry (e.g., blood plaque) for simulation. To evaluate the performance of our methods, we performed several experiments. The first experiment simulates the blood flow in an elastic cylinder-like vessel. For comparison, we use Finite Element Methods to solve the same problem using different grid resolutions and time steps. The results revealed that our approach can produce comparable results to FEM in high-resolution grids while consuming notably less time. Furthermore, we perform the blood flow simulation on a vessel with a plaque attained further to illustrate the robustness of our methods on complex geometry. In addition, we performed self-ablation on networks with different depths and widths to show how we should choose the architecture for the neural networks for our task. Since our methods are easy to parallelize, an evaluation of how our method can be accelerated by using multiple GPUs is performed. Last but not least, the mechanical analysis of how different sizes of plaque might affect the blood flow dynamics is done by using our methods.
@@
    We developed a neural solver based on deep learning techniques for fluid-structure interaction problems, which simulate the blood flow in deformable vessels. To the best of our knowledge, we are the first to use PINN method to solve an FSI problem.
@@
    Our method is mesh-free, removing the need for careful discretization and meshing of the computation domain. This feature makes it easy and robust to solve the simulation problems in different geometries.
@@
    The continuous nature of our method ensures that the solutions can be computed for any point in the domain. This advances the conventional grid-based methods, which will introduce extra errors by interpolating for points that are not nodes in the grid.
@@
    Benefiting from the well-developed deep learning framework packages, our method is easy to deploy on GPU and simple to parallelize. This greatly saves the computation time needed for solving the simulation problems.
@@ 2 Problem Formulation
Figure 1: The illustration for the Arbitrary Lagrangian Euler form. Left: reference frame; Right: current frame.
@@
To accurately represent the deformation and simulate the flow, it is crucial to properly model the movement of the computational domain, as well as the fluid velocity and pressure. As illustrated in Figure 1, we start with the reference frame $\Omega_{0}$, which serves as the parameterization for the entire domain required for our computations. Dividing this region into two parts, we have the fluid domain $\Omega_{0}^{\text{f}}$, through which the flow occurs, and the structural domain $\Omega_{0}^{\text{s}}$, representing the vessel wall. The interface between the fluid and structural domains is denoted as $\Gamma_{0}^{\text{w}}$. It’s important to note that most of the interaction takes place at this interface. The inlet boundary is labeled as $\Gamma_{0}^{\text{in}}$ and the outlet boundary is as $\Gamma_{0}^{\text{out}}$. Each point in the domain of the reference frame is denoted as the problem coordinates $\boldsymbol{x}_{0}=(x_{0},y_{0},z_{0})^{\intercal}$, where the bold symbols are used throughout this paper to signify matrices or column vectors. As the domain continues to move, we use $\Omega_{t}$ to represent the current frame at time $t$. The symbols $\Omega_{t}^{\text{f}}$, $\Omega_{t}^{\text{s}}$, $\Gamma_{0}^{\text{w}}$, $\Gamma_{0}^{\text{in}}$, and $\Gamma_{0}^{\text{out}}$ denote the fluid domain, the structural domain, the interface, the inlet, and the outlet, respectively. The coordinates of points in the current frame are referred to as the physical coordinates $(x_{t},y_{t},z_{t})^{\intercal}$.
Following Quarteroni et al. [26], we introduce the ALE mapping defined as:
$\mathcal{A}_{t}:\Omega_{0}\rightarrow\Omega_{t},\quad\boldsymbol{x}_{0}\rightarrow\boldsymbol{x}_{t}=\boldsymbol{x}_{0}+\boldsymbol{d}(\boldsymbol{x}_{0},t)=\mathcal{A}_{t}(\boldsymbol{x}_{0})\quad\text{ in }\Omega_{t}^{\text{f}},$ (1)
where the displacement function $\boldsymbol{d}(\cdot)$ characterizes the domain movement within the reference frame, and $\tilde{\boldsymbol{w}}=\frac{\partial\boldsymbol{d}}{\partial t}$ denotes the domain velocity within $\Omega_{0}$. In the current frame $\Omega_{t}$, the domain velocity is defined as $\boldsymbol{w}=\tilde{\boldsymbol{w}}\circ\mathcal{A}_{t}^{-1}$.
Suppose no external force is applied to the flow, the incompressible unsteady Navier-Stokes equation which discribes the movement of a viscous fluid in a moving domain is:
$\rho_{\text{f}}\frac{D}{Dt}\boldsymbol{u}+\rho_{\text{f}}[(\boldsymbol{u}-\boldsymbol{w})\cdot\nabla]\boldsymbol{u}-\nabla\cdot\boldsymbol{\sigma}_{\text{f}}=\boldsymbol{0}\\ \nabla\cdot\boldsymbol{u}=0\end{aligned}\quad\text{ in }\Omega_{t}^{\text{f}},\ t\in[0,T],$ (2)
@@
Figure 2: The coordinate system illustration. A 3D geometry can be represented using a Cartesian coordinate system as $(x,y,z)$ (left). A cylindrical model of the vessel geometry may be described by using a curvilinear cylindrical coordinate system $r=R(\theta,z;t)$ with the corresponding base unit vectors $\mathbf{e}_{r}$, $\mathbf{e}_{\theta}$ and $\mathbf{e}_{z}$, where $\mathbf{e}_{z}$ is aligned with the axis of the artery (right).
Assuming the vessel has a cylinder-like geometry at time 00, as illustrated in Figure 2, the vessel wall can be rewritten as:
$\Gamma_{0}^{\text{w}}=\left\{(r,\theta,z):r=R_{0}(\theta,z),\theta\in[0,2\pi),z\in[0,L]\right\},$ (3)
where $R_{0}(\theta,z)=r_{0}$, representing the initial radius of the cylindrical vessel geometry, and $L$ denotes the length of the arterial element under consideration. In our cylindrical coordinate system $(r,\theta,z)$, we assume that displacements are mainly in the radial direction and other components are negligible. Thus, we can simplify the displacement $\boldsymbol{d}$ to have only the radial component $\boldsymbol{\eta}$, such that:
$\boldsymbol{d}=\eta\boldsymbol{e}_{r},$ (4)
Here, $\boldsymbol{e}_{r}$ is the unit vector along the $r$-axis.
Given these settings, we rewrite the current frame $\Gamma_{t}^{w}$ as:
$\Gamma_{t}^{\text{w}}=\{(r,\theta,z):r=R(\theta,z;t),\theta\in[0,2\pi),z\in[0,L]\},$ (5)
@@
Based on linear elasticity and Newton’s second law of momentum, we equate them and obtain the stress continuity equation:
$\rho_{w}h_{0}R_{0}\frac{\partial^{2}\eta}{\partial t^{2}}+\frac{Eh_{0}}{1-\xi^{2}}\frac{\eta}{R}=-(2\mu\boldsymbol{D}(\boldsymbol{u})\cdot\boldsymbol{n})\cdot\boldsymbol{e}_{r}gR_{0}+pR\quad\text{ on }\Gamma_{0}^{\text{w}},$ (6)
where $g=\frac{R}{R_{0}}\sqrt{1+\left(\frac{1}{R}\frac{\partial R}{\partial\theta}\right)^{2}+\left(\frac{\partial R}{\partial z}\right)^{2}}$, $\rho_{\text{w}}$ is the vessel wall density, $h_{0}$ is the vessel wall thickness, $\xi$ is the Poisson ratio and $E$ is the Young modulus. ??? Refer to a reference for the derivation of this equation. ??? In this equation, the righthand side is the force exerted by the fluid, while the lefthand side is the internal force of the vessel wall structure. The first term on the left is the acceleration times mass by Newton’s second law of motion, and the second term is the linear elasticity.
Rewrite (6), we obtain the independent ring model for the stress continuity equation:
$\frac{\partial^{2}\eta}{\partial t^{2}}+b\eta=H\quad\text{ in }\Gamma_{0}^{\text{w}},$ (7)
@@
The structural domain in the reference frame is then denoted as:
$\Omega_{0}^{s}=\left\{(r,\theta,z):r\in[r_{0},r_{0}+h],\theta\in[0,2\pi),z\in[0,L]\right\}.$ (8)
We assume that the magnitude of the displacement remains constant along the radial direction. Then, the displacement $\boldsymbol{d}$ in the structure domain on the reference frame is expressed as
$\boldsymbol{d}(r_{\text{s}},\theta,z,t)=\eta(\theta,z;t)\boldsymbol{e}_{r}+r_{\text{s}}\boldsymbol{e}_{r}\quad\text{ for }(r_{\text{s}},\theta,z)\in\Omega_{0}^{s},t\in[0,T].$ (9)
In addition to modeling, we need to define the initial and boundary conditions. Given the variety of specific problem settings, we provide a general representation and define them specifically in the experimental section. For fluid problem, the boundary condition is given by:
$F^{\text{f}}_{\text{bdr}}(\boldsymbol{u},P)=0\quad\text{ on }\Gamma_{t},t\in[0,T],$ (10)
where $\Gamma_{t}=\Gamma^{\text{in}}_{t}\cup\Gamma^{\text{out}}_{t}\cup\Gamma^{\text{w}}_{t}$. The initial condition can be written as :
$F^{\text{f}}_{\text{init}}(\boldsymbol{u},P)=0\quad\text{ in }\Omega_{0}^{\text{f}},t=0.$ (11)
@@
$F^{\text{s}}_{\text{bdr}}(\boldsymbol{d})=0\quad\text{ on }\partial\Gamma^{\text{w}}_{0},t\in[0,T],$ (12)
@@
$F^{\text{s}}_{\text{init}}(\boldsymbol{d})=0\quad\text{ on }\Gamma^{\text{w}}_{0},t=0.$ (13)
@@
$\Delta\boldsymbol{d}=\boldsymbol{0}\quad\text{ in }\Omega_{0}^{\text{f}},t\in[0,T].$ (14)
@@ 3.1 Fully connected neural network
The most fundamental and widely utilized type of neural network in applied mathematics is the fully connected neural network (FNN). Let us denote the set of positive integers as $\mathbb{N}^{+}$. Mathematically, we define a simple nonlinear function, labeled as $h_{\ell}:\mathbb{R}^{M_{\ell-1}}\rightarrow\mathbb{R}^{M_{\ell}}$, where $M_{\ell}\in\mathbb{N}^{+}$ for $\ell=0,1,\ldots,L$. The expression for $h_{\ell}$ can be represented as
$h_{\ell}\left(\boldsymbol{x}_{\ell}\right):=\phi\left(\boldsymbol{W}_{\ell}\boldsymbol{x}_{\ell}+\boldsymbol{b}_{\ell}\right),$ (15)
@@
$\hat{\boldsymbol{x}}=(x_{1},x_{2},\cdots,x_{d-1},t)^{\intercal},$ (16)
@@
$N(\boldsymbol{x},t;{\boldsymbol{\theta}})=h_{L}\circ h_{L-1}\circ h_{L-2}\circ\cdots\circ h_{1}(\hat{\boldsymbol{x}})\quad\text{ for }\hat{\boldsymbol{x}}\in\mathbb{R}^{d},$ (17)
@@
To ensure the smoothness of the solutions approximated by neural networks, the inclusion of the Sigmoid activation function in some of our networks is essential. However, based on our experiments, we observed that utilizing only the Sigmoid activation function in the network can result in slow optimization and convergence. This issue may stem from the problem of gradient explosion associated with the Sigmoid [28, 16].
In response to this challenge, we have thoughtfully incorporated a combination of alternating activation functions into our network architecture. Specifically, after the first layer, which is activated by the Sigmoid function, in this work, we set the activation function for odd-numbered layers as Sigmoid while setting that in the even-numbered layer as ReLU. Besides, it’s important to note that the layer before the final output layer does not incorporate any activation function so as to ensure the correct scaling. This deliberate choice of activation functions strikes a balance between ensuring sufficiently continuous derivatives, which are crucial for precise fluid dynamics modeling, and the fast convergence offered by the Relu activation.

### v4

@@ 1 Introduction
To address these challenges, we propose a neural solver for the Navier-Stokes equation in deformable vessels using Physics-Informed Neural Networks (PINNs)[30]. In our methods, we take some networks to produce displacement, pressure and velocity of our problem. To acquire those networks, we first convert the partial differential equations of the formulated physical model into residual forms and use them as loss functions. Next, the automatic differentiation is used to represent all the differential operators. Then, we train the networks to minimize those losses. With those losses getting approximately zero, the networks can be an approximation of the functions for displacement, pressure, and velocity that are used to solve the modeled problem.
Benefit from the PINNs method, the proposed method is a mesh-free approach, eliminating the need for discretization and meshing for the computational domain. This feature makes it highly efficient and straightforward to add other objects into the geometry (e.g., blood plaque) for simulation. To evaluate the performance of our methods, we performed several experiments. The first experiment simulates the blood flow in an elastic cylinder-like vessel. For comparison, we use Finite Element Methods to solve the same problem using different grid resolutions and time steps. The results revealed that our approach can produce comparable results to FEM in high-resolution grids while consuming notably less time. Furthermore, we perform the blood flow simulation on a vessel with a plaque attained further to illustrate the robustness of our methods on complex geometry. In addition, we performed self-ablation on networks with different depths and widths to show how we should choose the architecture for the neural networks for our task. Since our methods are easy to parallelize, an evaluation of how our method can be accelerated by using multiple GPUs is performed. Last but not least, the mechanical analysis of how different sizes of plaque might affect the blood flow dynamics is done by using our methods.
@@
    We incorporated the Navier-Stokes equation in the Arbitrary Lagrangian-Eulerian (ALE) form into the Physics-Informed Neural Network to tackle fluid-structure interaction problems in a meshless approach.
@@
    We designed a sequential-alternative training approach by conceptualizing our problem as a multi-objective optimization task encompassing fluid dynamics and deformation. This approach allows for stable training and better convergence by focusing on optimizing subsets of losses sequentially.
@@
    We introduced a novel architecture featuring separate networks for predicting velocity and pressure. This design choice aligns with the distinct operators governing these variables in the Navier-Stokes equation, thereby enabling effective capture of their unique characteristics and complexities.
@@
    We proposed a new architecture utilizing alternative activation functions by ReLU and Sigmoid. This modification results in faster training and improved convergence dynamics.
@@ 2 Problem Formulation
Figure 1: The illustration for the Arbitrary Lagrangian Euler form. Left: reference configuration; Right: current configuration.
@@
To accurately represent the deformation and simulate the flow, it is crucial to properly model the movement of the computational domain, as well as the fluid velocity and pressure. As illustrated in Figure 1, we start with the reference configuration $\Omega_{0}$, which serves as the parameterization for the entire domain required for our computations. Dividing this region into two parts, we have the fluid domain $\Omega_{0}^{\text{f}}$, through which the flow occurs, and the structural domain $\Omega_{0}^{\text{s}}$, representing the vessel wall. The interface between the fluid and structural domains is denoted as $\Gamma_{0}^{\text{w}}$. It’s important to note that most of the interaction takes place at this interface. The inlet boundary is labeled as $\Gamma_{0}^{\text{in}}$ and the outlet boundary is as $\Gamma_{0}^{\text{out}}$. Each point in the domain of the reference configuration is denoted as the problem coordinates $\boldsymbol{x}_{0}=(x_{0},y_{0},z_{0})^{\intercal}$, where the bold symbols are used throughout this paper to signify matrices or column vectors. As the domain continues to move, we use $\Omega_{t}$ to represent the current configuration at time $t$. The symbols $\Omega_{t}^{\text{f}}$, $\Omega_{t}^{\text{s}}$, $\Gamma_{0}^{\text{w}}$, $\Gamma_{0}^{\text{in}}$, and $\Gamma_{0}^{\text{out}}$ denote the fluid domain, the structural domain, the interface, the inlet, and the outlet, respectively. The coordinates of points in the current configuration are referred to as the physical coordinates $(x_{t},y_{t},z_{t})^{\intercal}$.
Following Quarteroni et al. [29, P. 73, Sec. 18], we introduce the ALE mapping defined as:
(1) $\mathcal{A}_{t}:\Omega_{0}\rightarrow\Omega_{t},\quad\boldsymbol{x}_{0}\rightarrow\boldsymbol{x}_{t}=\boldsymbol{x}_{0}+\boldsymbol{d}(\boldsymbol{x}_{0},t)=\mathcal{A}_{t}(\boldsymbol{x}_{0})\quad\text{ in }\Omega_{t}^{\text{f}},$
where the displacement function $\boldsymbol{d}(\cdot)$ characterizes the domain movement within the reference configuration, and $\tilde{\boldsymbol{w}}=\frac{\partial\boldsymbol{d}}{\partial t}$ denotes the domain velocity within $\Omega_{0}$. In the current domain $\Omega_{t}$, the domain velocity is defined as $\boldsymbol{w}=\tilde{\boldsymbol{w}}\circ\mathcal{A}_{t}^{-1}$.
Suppose no external force is applied to the flow, the incompressible unsteady Navier-Stokes equation which describes the movement of a viscous fluid in a moving domain [29, P. 74, Sec. 18] is:
(2) $\rho_{\text{f}}\frac{D}{Dt}\boldsymbol{u}+\rho_{\text{f}}[(\boldsymbol{u}-\boldsymbol{w})\cdot\nabla]\boldsymbol{u}-\nabla\cdot\boldsymbol{\sigma}_{\text{f}}=\boldsymbol{0}\\ \nabla\cdot\boldsymbol{u}=0\end{aligned}\quad\text{ in }\Omega_{t}^{\text{f}},\ t\in[0,T],$
@@
Assuming the vessel has a cylinder-like geometry at time 00, which is also the reference domain, the vessel wall [29, P. 60, Sec. 16] can be rewritten as:
(3) $\Gamma_{0}^{\text{w}}=\left\{(r,\theta,z):r=R_{0}(\theta,z),\theta\in[0,2\pi),z\in[0,L]\right\},$
where $R_{0}(\theta,z)$ is the radius for the vessel lumen at $(\theta,z)$. Assume the vessel geometry in reference configuration to be a cylinder-like geometry, we have $R_{0}(\theta,z)=r_{0}$ with $r_{0}\in\mathbb{R}$ is a constant, representing the initial radius of the cylindrical vessel geometry, and $L$ denotes the length of the arterial element under consideration. In our cylindrical coordinate system $(r,\theta,z)$, we assume that displacements are mainly in the radial direction and other components are negligible. Thus, we can simplify the displacement $\boldsymbol{d}$ to have only the radial component ${\eta}$, such that:
(4) $\boldsymbol{d}(r,\theta,z;t)=\eta(\theta,z;t)\boldsymbol{e}_{r},$
Here, $\eta(\theta,z;t)$ is a function for the displacement at the vessel wall, and $\boldsymbol{e}_{r}$ is the unit vector along the radial direction.
Given these settings, we can write the domain for the vessel wall $\Gamma_{t}^{w}$ at time $t$ as:
(5) $\Gamma_{t}^{\text{w}}=\{(r,\theta,z):r=R(\theta,z;t),\theta\in[0,2\pi),z\in[0,L]\},$
@@
The same as [29, P.64, Sec. 16], based on linear elasticity and Newton’s second law of momentum, we equate the internal force of the structure to the force exerted from the fluid, obtaining the below stress continuity equation on the vessel wall:
(6) $\rho_{w}h_{0}R_{0}\frac{\partial^{2}\eta}{\partial t^{2}}+\frac{Eh_{0}}{1-\xi^{2}}\frac{\eta}{R}=-(2\mu\boldsymbol{D}(\boldsymbol{u})\cdot\boldsymbol{n})\cdot\boldsymbol{e}_{r}gR_{0}+PR\quad\text{ on }\Gamma_{0}^{\text{w}},$
where $g=\frac{R}{R_{0}}\sqrt{1+\left(\frac{1}{R}\frac{\partial R}{\partial\theta}\right)^{2}+\left(\frac{\partial R}{\partial z}\right)^{2}}$, $\boldsymbol{D}(\boldsymbol{u})=\frac{\nabla\boldsymbol{u}+(\nabla\boldsymbol{u})^{\intercal}}{2}$, $\rho_{\text{w}}$ is the vessel wall density, $h_{0}$ is the vessel wall thickness, $\xi$ is the Poisson ratio and $E$ is the Young modulus[29].
Regard $\boldsymbol{u}$ as a known, divide (6) by $\rho_{w}h_{0}R_{0}$, we obtain the independent ring model for the stress continuity equation [29, Eq. (16.12), P. 60, Sec. 16]:
(7) $\frac{\partial^{2}\eta}{\partial t^{2}}+b\eta=H\quad\text{ in }\Gamma_{0}^{\text{w}},$
@@
The structural domain in the reference configuration is then denoted as:
(8) $\Omega_{0}^{s}=\left\{(r,\theta,z):r\in[r_{0},r_{0}+h_{0}],\theta\in[0,2\pi),z\in[0,L]\right\}.$
We assume that the magnitude of the displacement remains constant along the radial direction. Then, the displacement $\boldsymbol{d}$ in the structure domain on the reference configuration is expressed as
(9) $\boldsymbol{d}(r,\theta,z;t)=\eta(\theta,z;t)\boldsymbol{e}_{r}+r\boldsymbol{e}_{r}\quad\text{ for }(r,\theta,z)\in\Omega_{0}^{s},t\in[0,T].$
In addition to modeling, we need to define the initial and boundary conditions. Given the variety of specific problem settings, we provide a general representation and define them specifically in the experimental section.
For Navier Stokes equation, which we can refer to as the fluid problem, the boundary condition is given by:
(10) $F^{\text{f}}_{\text{bdr}}(\boldsymbol{u},P)=0\quad\text{ on }\Gamma_{t},t\in[0,T],$
where $\Gamma_{t}=\Gamma^{\text{in}}_{t}\cup\Gamma^{\text{out}}_{t}\cup\Gamma^{\text{w}}_{t}$. Note that (10) includes the inlet boundary condition, outlet boundary condition and the interface condition, which will be explicitly given later in Section.4.
The initial condition can be written as :
(11) $F^{\text{f}}_{\text{init}}(\boldsymbol{u},P)=0\quad\text{ in }\Omega_{0}^{\text{f}},t=0.$
@@
(12) $F^{\text{s}}_{\text{bdr}}(\boldsymbol{d})=0\quad\text{ on }\partial\Gamma^{\text{w}}_{0},t\in[0,T],$
@@
(13) $F^{\text{s}}_{\text{init}}(\boldsymbol{d})=0\quad\text{ on }\Gamma^{\text{w}}_{0},t=0.$
@@
(14) $\Delta\boldsymbol{d}=\boldsymbol{0}\quad\text{ in }\Omega_{0}^{\text{f}},t\in[0,T].$
@@ 3.1 Fully connected neural network
Denoting the set of positive integers as $\mathbb{N}^{+}$, we define a simple nonlinear function, labeled as $h_{\ell}:\mathbb{R}^{M_{\ell-1}}\rightarrow\mathbb{R}^{M_{\ell}}$, where $M_{\ell}\in\mathbb{N}^{+}$ for $\ell=0,1,\ldots,L$. The expression for $h_{\ell}$ can be represented as
(15) $h_{\ell}\left(\boldsymbol{x}_{\ell}\right):=\phi\left(\boldsymbol{W}_{\ell}\boldsymbol{x}_{\ell}+\boldsymbol{b}_{\ell}\right),$
@@
(16) $\hat{\boldsymbol{x}}=(x_{1},x_{2},\cdots,x_{d-1},t)^{\intercal},$
@@
(17) $N(\boldsymbol{x},t;{\boldsymbol{\theta}})=h_{L}\circ h_{L-1}\circ h_{L-2}\circ\cdots\circ h_{1}(\hat{\boldsymbol{x}})\quad\text{ for }\hat{\boldsymbol{x}}\in\mathbb{R}^{d},$
@@
To ensure the smoothness of the solutions approximated by neural networks, the inclusion of the Sigmoid activation function in some of our networks is essential. However, based on our experiments, we observed that utilizing only the Sigmoid activation function in the network can result in slow optimization and convergence. This issue may stem from the problem of vanishing gradient associated with the Sigmoid [31, 16].
In response to this challenge, we have thoughtfully incorporated a combination of alternating activation functions into our network architecture. Specifically, after the first layer, which is activated by the Sigmoid function, in this work, we set the activation function for odd-numbered layers as Sigmoid while setting that in the even-numbered layer as ReLU. Besides, it’s important to note that the layer before the final output layer does not incorporate any activation function so as to ensure the correct scaling. This deliberate choice of activation functions strikes a balance between ensuring sufficiently continuous derivatives from Sigmoid, which are crucial for precise fluid dynamics modeling[19], and the fast convergence offered by the ReLU activation[21].